# Neural Green’s Operators for Parametric Partial Differential Equations

### v4

@@ Abstract
This work introduces a paradigm for constructing parametric neural operators that are derived from finite-dimensional representations of Green’s operators, with learnable Green’s functions, for linear partial differential equations (PDEs). We refer to such neural operators as Neural Green’s Operators (NGOs). Our construction of NGOs preserves the linear action of Green’s operators on the inhomogeneity fields, while approximating the nonlinear dependence of the Green’s function on the coefficients of the PDE using neural networks that take weighted averages of such coefficients as input. This construction reduces the complexity of the problem from learning the entire solution operator and its dependence on all parameters to only learning the Green’s function and its dependence on the PDE coefficients. Moreover, taking weighted averages, rather than point samples, of input functions decouples the network size from the number of sampling points, enabling efficient resolution of multiple scales in the input fields. Furthermore, we show that our explicit representation of Green’s functions enables the embedding of desirable mathematical attributes in our NGO architectures, such as symmetry, spectral, and conservation properties. Through numerical benchmarks on canonical PDEs, we demonstrate that NGOs achieve comparable or superior accuracy to deep operator networks, variationally mimetic operator networks, and Fourier neural operators with similar parameter counts, while generalizing significantly better when tested on out-of-distribution data. For time-dependent PDEs, we show that NGOs can produce pointwise-accurate dynamics in an auto-regressive manner when trained on a single time step. Finally, we show that we can leverage the explicit representation of Green’s functions returned by NGOs to construct effective matrix preconditioners that accelerate iterative solvers for PDEs.
@@ 1 Introduction
In this paper, we consider the problem of inferring the solution operator, associated with a parametric family of partial differential equations (PDEs), from data. Generally, the task of approximating such solution operators plays a fundamental role in the development of computational methods for solving the underlying PDEs [1, 2] that appear in various areas of science and engineering. Typical examples include computational fluid dynamics [3], fluid-structure interaction [4] and solid mechanics [5]. However, finding analytical expressions of such solution operators is generally not feasible, and numerical approximations of such operators can be computationally prohibitive due to their high-dimensional and nonlinear nature [1]. Recent advances in operator learning techniques [6] address this challenge by using data to approximate solution operators to PDEs. Typically, the data comprises instantiations of solutions to the PDE and the corresponding PDE input functions, such as boundary conditions, PDE coefficients and external forcings. The inference problem is to find the operator that maps the PDE parameters to the solution. If such a mapping is inferred, we are able to evaluate it and recover solutions of the PDE for different input parameters.
@@
By virtue of their generalization of standard neural networks [7] to mappings between function spaces instead of between discrete vector spaces, neural operators (NOs) [6] provide a suitable representation of PDE solution maps. Representation of operators by NOs can be roughly divided into two categories. The first approach expands the output function in terms of the dot product between two vectors that are returned from two sub-networks, the so-called branch and trunk networks, which accept as arguments samples of the input functions and spatial coordinates, respectively. The second approach adds an integral operator to the discrete affine transformation in a standard neural network which allows application of the NO to varying resolutions of the input and output functions. The former construction is due to work in [8] and was extended to Deep Operator Networks (DeepONets) in [9], while the latter construction was developed for different choices of the nonlocal kernel functions such as the Fourier [10], multipole [11] and graph [12], neural operators (FNOs, MNOs and GNOs, respectively). Both such approaches received considerable research attention, and significant work has been devoted to their extensions and improvements [6]. Moreover, both approaches have been successfully applied to approximating solutions to PDEs in various fields, such as PDEs for fluid mechanics [13], fluid-structure-interaction [14], solid mechanics [15] and quantum mechanics [16]. While the errors incurred by such NOs generalize perform well to unseen PDE parameters within the training distribution, extrapolation beyond that distribution may become unreliable [17]. The lack of reliability in extrapolation may preclude the adoption of NOs, particularly when training data may be limited, e.g. in applications where the PDE parameters vary across length and time scales while training data may only cover a subset of those scales. Moreover, such generalization errors may be compounded when such NOs are used in an auto-regressive manner, such as predicting long-time behavior when trained on short trajectories [18, 19, 20]. Furthermore, standard application of such NOs [21, 9, 10] require point samples of the PDE parameters as input which scales unfavorably with the number of trainable parameters. This may result in inefficient models for multi-scale systems where a dense sampling of input functions is required.
@@
    NGOs can produce pointwise-accurate dynamics of time-dependent PDEs in an auto-regressive manner over arbitrarily large numbers of time steps, when trained on a single time step only.
@@
The remainder of this paper is organized as follows: Section 2 introduces Green’s operators for parametric PDEs and corresponding finite-dimensional representations of the Green’s function. In Section 3, we introduce variants of NGO architectures for different modeling scenarios. In Section 4, we describe a canonical elliptic PDE problem and the corresponding data generation and training procedure. Section 4.3 provides results on the model accuracy, the effect of quadrature, and computational efficiency. Additionally, we demonstrate that the learned Green’s functions can be used to construct effective preconditioners. In Section 5, we test our NGO framework on some non-elliptic PDE problems. We end with a concluding discussion in Section 6.
@@ 3.1.3 Data NGO
Table 1 compares some key properties of the different models. NGOs are the only tested models for which (1) the input data can be given on an arbitrary set of points (provided an accurate quadrature rule is available), (2) the output can be evaluated anywhere, and (3) the linear structure of the PDE is preserved.
@@
Model Input samples Output samples Linearity
DeepONet Fixed samples Continuous No
VarMiON Fixed samples Continuous Yes
FNO Uniform grid Same as input No
UNet Fixed uniform grid Same as input No
CNO Uniform grid Uniform grid No
NGOs Continuous Continuous Yes
@@ 4 Test Problem: Steady diffusion
In this section, we compare the performance of the NGO with other models in literature, including DeepONets, VarMiONs, FNOs, CNOs, and U-Nets. The test equation is a steady diffusion equation on a unit square, with varying diffusion coefficient $\theta$. The boundary conditions are Neumann boundary conditions on the top and bottom boundaries, and Dirichlet on the left and right boundaries. The equations are
$\begin{cases}\hfill-\nabla\cdot\left(\theta\nabla u\right)=f&\text{on }\Omega=(0,L)^{2},\\ \hfill\theta\nabla u\cdot\mathbf{n}=\eta&\text{on }\Gamma_{\mathrm{N}}=(0,L)\times\left\{0,L\right\},\\ \hfill\theta u=g&\text{on }\Gamma_{\mathrm{D}}=\left\{0,L\right\}\times(0,L),\end{cases}$ (4.1)
@@
Figure 4: The square domain $\Omega$, with Neumann boundaries $\Gamma_{\mathrm{N}}$ on the top and bottom, and Dirichlet boundaries $\Gamma_{\mathrm{D}}$ on the left and right.
@@ 4.3.1 Generalization Errors of Different NOs
The errors are given in terms of mean and standard deviation of the relative $L^{2}$ error. We define the relative $L^{2}$ error as $e_{i}=\frac{\|\hat{u}^{(i)}-u^{(i)}\|_{L^{2}(\Omega)}}{\|u^{(i)}\|_{L^{2}(\Omega)}}$, where $u^{(i)},i=1,\dots,N$ are the true solutions and $\hat{u}^{(i)}$ are the model predictions.
@@ 4.3.2 Generalization Errors Across Length Scales
Figure 8: Relative $L^{2}$ test error versus test data length scale $\lambda/L$ for a bare NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO. All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO, also added as reference) use the same $10\times 10$ cubic B-spline basis, and are thus limited by the same $L^{2}$ projection error lower bound, indicated in black. Points and error bars are, respectively, averages and 95% confidence intervals on datasets of 1000 manufactured steady diffusion solutions, generated in the same way as dataset C.
@@
Figure 9: The input structure of an NGO compared to input structure of neural operators from literature like the DeepONet, FNO and VarMiON. Typically, the learnable components of canonical neural operators take as input the input function $\theta$ sampled on a quadrature grid $\mathbf{x}_{q}$, so the size of the input vector is determined by the number of quadrature points $Q$. On the other hand, the NGO system net takes as input inner products of the input function with a basis, which are computed using a quadrature rule with weights $w_{q}$ and points $\mathbf{x}_{q}$. The size of the NGO input is thereby determined by the size of the basis NN, and independent of the number of used quadrature points $Q$. Therefore, for the same NN input size (in the example illustration, 5 numbers), an NGO can use (arbitrarily) finer quadrature, resulting in higher accuracy on problems that involve fine scale data, where dense quadrature is required.
@@ 4.3.6 NGO-based Preconditioners
Recently, neural operators have gained interest as methods to speed up the numerical solution of PDEs, using discretizations such as finite differences or finite elements. There are multiple ways to achieve this acceleration. One method, used by Zhang et al. [24], Hu et al. [25], and Ackmann et al. [26] among others, constructs a fixed-point iteration scheme for linear systems (derived from PDEs), that alternates between a “standard” smoothing step, such as Jacobi or Gauss-Seidel, and an application of the respective neural operator. Others, including Kopaničáková et al. [27] and Xiang et al. [28], use the learned operator as a preconditioner directly, in combination with a linear solver such as Flexible GMRES [29] to allow non-linear operators as preconditioners for linear systems. However, neither approach is generally applicable to PDEs as they require specific choices for the solver or iteration scheme used.
An advantage of NGOs is that the fact that they learn an approximation to the Green’s function facilitates the construction of preconditioners for numerical discretizations of the same PDE. In this section, we give an example of using a trained NGO to construct a preconditioner for a finite difference discretization for the diffusion equation. Specifically, a given discretization of a PDE yields a linear system with a matrix $\mathbf{C}[\theta]$, for which we can construct a right preconditioner $\mathbf{M}[\theta]$ from the NGO, such that the condition number of $\mathbf{P}=\mathbf{CM}^{-1}$ is much smaller than that of $\mathbf{C}$.The approach taken here loosely follows that of Pasetto et al. [30], in which a preconditioner for the time-varying diffusion is constructed from a reduced-order model. There, the discretised differential operator is restricted to a reduced bases, and the inverse of this coarse-scale operator is taken as a starting point for a preconditioner. In our approach, both the reduced basis and the approximate solution operator on this basis are taken from the trained NGO.
For this example, we take the steady diffusion problems from the testing data set, formulate a finite difference discretization of the PDE, and use the data NGO trained in the previous section to construct a preconditioner. For the finite difference discretization, we assume a uniform grid (i.e. $\Delta x=\Delta y\equiv h$) and use a five-point difference scheme to approximate the differential operator:
@@
where $\theta_{i,j}=\theta(x_{i},y_{j})$ and $u_{i,j}$ is the finite-difference approximation to $u(x_{i},x_{j})$. We take $h=\frac{1}{99}$, which produces a discretisation with $100$ points in the $y$-direction and $98$ points in the $x$-direction (the Dirichlet boundary conditions imply that no boundary points need to be included). This produces a linear system $\mathbf{C}\mathbf{u}=\mathbf{b}$.
The linear system has $N=9800$ unknowns, which is much larger than the $n=144$ basis functions of the NGO. This means that we cannot directly apply the NGO as a preconditioner, since the NGO learns an operator of rank $144$ at most. Applying it as a preconditioner directly would result in a singular system, which does not have a unique solution.
Nevertheless, we can use the trained NGO to construct a preconditioner $\mathbf{M}$ such that $\kappa(\mathbf{CM}^{-1})\ll\kappa(\mathbf{C})$. Generally, discretized differential operators may produce ill-conditioned matrices since they amplify high-frequency components while dampening low-frequency components. As such, a preconditioning strategy is to construct a preconditioner $\mathbf{M}$ that amplifies the low frequencies, while preserving the high frequencies. To construct such a preconditioner, we first write the learned Green’s function in a singular value decomposition (SVD) form.
Note that since the NGOs for the diffusion problem are of Galerkin form, i.e. with $\psi_{m}=\phi_{m}$, the separate test basis $\psi_{m}$ is omitted here, and $\phi_{m}$ is used for both the test and trial bases.
@@
    For the given parameters $\theta$, the NGO learns an approximation to the Green’s operator as in (2.8):
    $\hat{G}[\theta](\mathbf{x},\mathbf{x}^{\prime})$ $\approx\sum_{m,n}\phi_{m}(\mathbf{x})\mathbf{A}_{mn}[\theta]\phi_{n}(\mathbf{x}^{\prime}).$ (4.17)
@@
    Now, we first orthonormalize the function basis: define the Gram matrix $\mathbf{G}_{mn}=\int_{\Omega}\psi_{m}(\mathbf{x})\phi_{n}(\mathbf{x})~d\mathbf{x}$, and take the Cholesky decomposition $\mathbf{G}=\mathbf{LL}^{\top}$. Then define $\tilde{\phi}_{m}(\mathbf{x})=\sum_{m}\mathbf{L}^{-1}_{mn}\phi_{n}(\mathbf{x})$, which gives us a representation of the learned Green’s function in terms of an orthonormal basis:
    $\hat{G}[\theta](\mathbf{x},\mathbf{x}^{\prime})$ $\approx\sum_{m,n}\tilde{\phi}_{m}(\mathbf{x})\left(\mathbf{L}^{\top}\mathbf{A}(\theta)\mathbf{L}\right)_{mn}\tilde{\phi}_{n}(\mathbf{x}^{\prime}).$ (4.18)

### v5

@@ Abstract
00footnotetext: These authors contributed equally to this work
This work introduces a paradigm for constructing parametric neural operators that are derived from finite-dimensional representations of Green’s operators for linear partial differential equations (PDEs). We refer to such neural operators as Neural Green’s Operators (NGOs). Our construction of NGOs preserves the linear action of Green’s operators on the inhomogeneity fields, while approximating the nonlinear dependence of the Green’s function on the coefficients of the PDE using neural networks that take weighted averages of such coefficients as input. This construction reduces the complexity of the problem from learning the entire solution operator and its dependence on all parameters to only learning the Green’s function and its dependence on the PDE coefficients. Moreover, taking weighted averages, rather than point samples, of input functions decouples the network size from the number of sampling points, enabling efficient resolution of multiple scales in the input fields. Furthermore, we show that our explicit representation of Green’s functions enables the embedding of desirable mathematical attributes in our NGO architectures, such as symmetry, spectral, and conservation properties. Through numerical benchmarks on canonical PDEs, we demonstrate that NGOs achieve comparable or superior accuracy to Deep Operator Networks, Variationally Mimetic Operator Networks, and Fourier Neural Operators with similar parameter counts, while generalizing significantly better when tested on out-of-distribution data. For parametric time-dependent PDEs, we show that NGOs that are trained on a single time step can produce pointwise-accurate dynamics in an auto-regressive manner over arbitrarily large numbers of time steps. For parametric nonlinear PDEs, we demonstrate that NGOs trained exclusively on solutions of corresponding linear problems can be embedded within iterative solvers to yield accurate solutions, provided a suitable initial guess is available. Finally, we show that we can leverage the explicit representation of Green’s functions returned by NGOs to construct effective matrix preconditioners that accelerate iterative solvers for PDEs.
@@ 1 Introduction
In this paper, we consider the problem of inferring the solution operator, associated with a parametric family of partial differential equations (PDEs), from data. Generally, the task of approximating such solution operators plays a fundamental role in the development of computational methods for solving the underlying PDEs [23, 24] that appear in various areas of science and engineering. Typical examples include computational fluid dynamics [2], fluid-structure interaction [8] and solid mechanics [14]. However, finding analytical expressions of such solution operators is generally not feasible, and numerical approximations of such operators can be computationally prohibitive due to their high-dimensional and nonlinear nature [23]. Recent advances in operator learning techniques [38, 53] address this challenge by using data to approximate solution operators to PDEs. Typically, the data comprises instantiations of solutions to the PDE and the corresponding PDE input functions, such as boundary conditions, PDE coefficients and external forcings. The inference problem is to find the operator that maps the PDE parameters to the solution. If such a mapping is inferred, we are able to evaluate it and recover solutions of the PDE for different input parameters.
@@
By virtue of their generalization of standard neural networks [20] to mappings between function spaces instead of between discrete vector spaces, neural operators (NOs) [38] provide a suitable representation of PDE solution maps. Representation of operators by NOs can be roughly divided into two categories. The first approach expands the output function in terms of the dot product between two vectors that are returned from two sub-networks, the so-called branch and trunk networks, which accept as arguments samples of the input functions and spatial coordinates, respectively. The second approach adds an integral operator to the discrete affine transformation in a standard neural network which allows application of the NO to varying resolutions of the input and output functions. The former construction is due to work in [10] and was extended to Deep Operator Networks (DeepONets) in [37], while the latter construction was developed for different choices of the nonlocal kernel functions such as the Fourier [34], multipole [35] and graph [30], neural operators (FNOs, MNOs and GNOs, respectively). Both such approaches received considerable research attention, and significant work has been devoted to their extensions and improvements [38]. Moreover, both approaches have been successfully applied to approximating solutions to PDEs in various fields, such as PDEs for fluid mechanics [29], fluid-structure-interaction [61], electromagnetism [4], solid mechanics [13] and quantum mechanics [19]. While such NOs generalize well to unseen PDE parameters within the training distribution, extrapolation beyond that distribution may become unreliable [63]. The lack of reliability in extrapolation may preclude the adoption of NOs, particularly when training data may be limited, e.g. in applications where the PDE parameters vary across length and time scales while training data may only cover a subset of those scales. Moreover, such generalization errors may be compounded when such NOs are used in an auto-regressive manner, such as predicting long-time behavior when trained on short trajectories [40, 39, 9]. Furthermore, standard application of such NOs [43, 37, 34] require point samples of the PDE parameters as input which scales unfavorably with the number of trainable parameters. This may result in inefficient models for multi-scale systems where a dense sampling of input functions is required. This shortcoming can be addressed by constructing neural operators that act on functions in a discretization-independent manner to improve the interpretability and accuracy of the model [5, 22, 6].
@@
    NGOs that are trained on a single time step can produce pointwise-accurate dynamics of time-dependent PDEs in an auto-regressive manner over arbitrarily large numbers of time steps.
@@
    NGOs trained exclusively on linearized parametric PDEs can be embedded within iterative nonlinear solvers to yield accurate solutions to associated nonlinear problems, provided a suitable initial guess is available.
  4. 4.
@@
1.1 Other Related Work
So far, our discussion has focused on neural operators for families of PDEs indexed by their coefficients, and the placement of neural Green’s operators in such a class of networks. Since NGOs learn solution operators to parametric PDEs by mimicking the mathematical construction of Green’s operators in the neural architecture, NGOs also learn the associated parametric Green’s functions. This generalization to the parametric setting distinguishes our work from prior research devoted to inferring Green’s functions [boullé2022data, 12, 15, 17, 44, 56, 25]. By learning the parametrization of the Green’s function with respect to the PDE coefficients, our learned Green’s operator can be used as a versatile numerical building block; as demonstrated in Sections 5.1 and 5.3, the trained neural Green’s operator can be embedded into iterative algorithms for time-dependent and nonlinear problems. For the problems considered in this work, such tasks would be infeasible with non-parametric Green’s functions. Furthermore, we show in Section 4.3.6 that the inferred parametric Green’s functions can be used to construct classical matrix preconditioners for parametric linear systems. Unlike hybrid neural approaches to preconditioning [3, 11, 31, 62, kopaničáková2025deeponet, 60], our learned matrix preconditioner can be applied algebraically to any Krylov subspace method. In the remainder of this section we review such related work, both to situate our contribution within the broader literature and to emphasize the distinctive capabilities enabled by our formulation.
Significant research work has been devoted to inferring an explicit representation of Green’s functions for PDEs, for example in [boullé2022data],[44], [56] and [25], Green’s functions of linear PDEs are represented by neural networks; in [15] autoencoders are used to infer a matrix representation of the linearization of a PDE operator, where the inverse of such a matrix represents Green’s function; in [17] a Green’s function to a target PDE is represented by a trunk network of a DeepONet [37], while the branch network of the DeepONet is used to approximate the auxiliary gradients of the Green’s function. In [12] network architectures that are inspired by Green’s function formulations are proposed to address challenges faced by some neural operators in accounting for boundary conditions. Despite these advances, the aforementioned approaches do not account for the (generally nonlinear) dependence of the Green’s function on the PDE coefficients and are therefore restricted to fixed PDE operators. An exception is presented in [45], where low-rank approximations of Green’s functions for fixed PDE coefficients are inferred from data, subsequently interpolating these representations across the PDE coefficient space. Nevertheless, this framework is restricted to self-adjoint operators, and its accuracy is critically contingent upon the regularity of the Green’s function and its spectral decomposition relative to the PDE coefficients. Therefore, a general and flexible framework for learning Green’s functions that vary parametrically with the PDE coefficients remains lacking. Addressing this gap is the central focus of the present work.
Apart from predicting approximate solutions to PDEs, neural operators can also be used to accelerate iterative methods for PDEs. For example, Ruelmann et al. [50] propose a neural network that maps system matrices to approximate inverses, thereby accelerating the solution of linear systems; however, this approach is limited to problems of a fixed size. Related efforts seek to learn mappings that output incomplete factorizations while preserving the sparsity pattern of the input matrix [häusner2024neural, 33, 32, 54], though these methods are, as presented there, restricted to matrices that are symmetric and positive definite. A distinct category of hybrid preconditioning, exemplified by the HINTS framework [21, 62, 26], integrates a solution-approximating neural operator directly into a fixed-point iteration. Similarly, neural operators have been embedded as preconditioners within Krylov-subspace solvers [60, 55, 31, 3]. A significant attribute of these existing works is their reliance on nonlinear models to approximate the action of what is fundamentally a linear operator, thus often meaning that Flexible GMRES (F-GMRES, [52]) is the only Krylov method for which the approach is applicable. In contrast, this work proposes algorithms for constructing matrix preconditioners that preserve the linear structure of classical preconditioners, thereby ensuring algebraic compatibility with the full range of standard Krylov subspace methods.
The remainder of this paper is organized as follows. Section 2 establishes the mathematical foundation of Green’s operators for parametric PDEs and introduces their corresponding finite-dimensional representations. In Section 3, we introduce the Neural Green’s Operator (NGO) and present architectural variants tailored to diverse modeling scenarios. Section 4 details the canonical elliptic PDE used as our primary benchmark, including the associated data generation and training protocols. Our core findings are presented in Section 4.3, where we evaluate model accuracy, computational efficiency, and the influence of critical hyperparameters such as quadrature schemes, basis function cardinality, and total parameter count. This section also demonstrates the utility of the learned Green’s functions in constructing robust matrix preconditioners. To assess the framework’s versatility, Section 5 explores the application of NGOs to non-elliptic, transient, and nonlinear problems. Finally, Section 6 provides a summary of our contributions and a discussion of future research directions.
@@ 3.1.3 Data NGO
Table 1 compares some key properties of the different models. NGOs are the only tested models for which (1) the input data can be given on an arbitrary set of points (provided an accurate quadrature rule is available), (2) the output can be evaluated anywhere, and (3) the linear structure of the PDE is preserved. Note that other machine learning methods related to Green’s functions [boullé2022data, 12, 15, 17, 44, 56], while perhaps conceptually similar to NGOs, are not included in the comparison as those models learn solutions to a single linear or nonlinear PDE, with no dependence on PDE parameters such as varying diffusion coefficients or advection fields. As such, these models cannot be trained to approximate the mappings that NGOs and other included neural operators learn.
@@
  [11] Model Input samples Output samples Linearity
[37] DeepONet Fixed samples Continuous No
[5] RINO Any quadrature rule Continuous No
[43] VarMiON Fixed samples Continuous Yes
[34] FNO Uniform grid Same as input No
[49] U-Net Fixed uniform grid Same as input No
[47] CNO Uniform grid Uniform grid No
  [11] NGOs Any quadrature rule Continuous Yes
@@ 4 Test Problem: Steady diffusion
In this section, we compare the performance of the NGO with other models in literature, including DeepONets, VarMiONs, FNOs, CNOs, and U-Nets. The test problem is the steady diffusion equation defined in a unit square and parameterized by a spatially varying diffusion coefficient $\theta(\mathbf{x})$, the Neumann boundary conditions $\eta(\mathbf{x})$ on the upper and lower boundaries and the Dirichlet boundary conditions $g(\mathbf{x})$ on the left and right boundaries; The equations are given by
$\begin{cases}\hfill-\nabla\cdot\left(\theta\nabla u\right)=f&\text{on }\Omega=(0,1)^{2},\\ \hfill\theta\nabla u\cdot\mathbf{n}=\eta&\text{on }\Gamma_{\mathrm{N}}=(0,1)\times\left\{0,1\right\},\\ \hfill\theta u=g&\text{on }\Gamma_{\mathrm{D}}=\left\{0,1\right\}\times(0,1),\end{cases}$ (4.1)
@@
Figure 4: The square domain $\Omega$, with Neumann boundaries $\Gamma_{\mathrm{N}}$ on the top and bottom, and Dirichlet boundaries $\Gamma_{\mathrm{D}}$ on the left and right. We used $L=1$.
@@ 4.1 NGO architecture
Remark 5.
The matrix $\mathbf{F}$ as given by 4.6 may not be coercive. Stabilization of such a formulation can be addressed using Nitsche’s method[7] which adds the terms
$F^{(\mathrm{s})}_{nm}=C_{\mathrm{s}}\int_{\Gamma_{\mathrm{D}}}\psi_{n}\theta\phi_{m}d\mathbf{x}^{\prime}$ (4.8)
and
$d^{(\mathrm{s})}_{n}=C_{\mathrm{s}}\int_{\Gamma_{\mathrm{D}}}\psi_{n}gd\mathbf{x}^{\prime}$ (4.9)
to the matrix $F_{nm}$ and right-hand-side vector $d_{n}$, respectively, where $C_{\mathrm{s}}>0$ is the Nitsche stabilization parameter. However, in our FEM tests in this section we observed that the stabilization is not needed. Therefore, we chose to not include the stabilization in the NGO formulation for the steady diffusion problem, because omitting it gave identical results, both for NGOs and FEM.
@@ 4.3.1 Generalization Errors of Different NOs
The errors are given in terms of mean and standard deviation of the relative $L^{2}$ error. We define the relative $L^{2}$ error as $e_{i}=\|\hat{u}^{(i)}-u^{(i)}\|_{L^{2}(\Omega)}/\|u^{(i)}\|_{L^{2}(\Omega)}$, where $u^{(i)},i=1,\dots,N$ are the true solutions and $\hat{u}^{(i)}$ are the model predictions.
@@
Remark 6.
We note that out of the models compared here, only the NGO and VarMiON architectures represent mappings that are linear in the source and boundary data of the PDE. In contrast, the DeepONet/RINO, FNO, CNO, and U-Net architectures can represent arbitrary (nonlinear) functions of $\theta$, $f$, $\eta$, and $g$. While this lack of structural constraint often results in higher out-of-distribution errors for these models, it also grants them the flexibility to represent a broader class of mappings. On the other hand, existing approaches that account for the linearity of the solution map in the source terms, such as [boullé2022data, 56], focus on a single instance of the PDE coefficients rather than a parametric family, and therefore are not directly comparable to NGOs. Finally, although we employ a parametric linear PDE as a benchmark, such problems are standard for evaluating even fully nonlinear neural operators [37, 34, 35, 47]. In Section 5.3 we demonstrate that NGOs can also be effective for nonlinear PDEs via fixed-point iteration.
@@ 4.3.2 Generalization Errors Across Length Scales
Figure 8: Relative $L^{2}$ test error versus test data length scale $\lambda/L$ (where we used $L=1$) for a standalone NN (here a U-Net), and the same NN, when used in a DeepONet, VarMiON, data NGO, data-free NGO and model NGO. All models have roughly $3\cdot 10^{4}$ trainable parameters, and all models (except the NN and FNO, also added as reference) use the same $10\times 10$ cubic B-spline basis, and are thus limited by the same $L^{2}$ projection error lower bound, indicated in black. Points and error bars are, respectively, averages and 95% confidence intervals on datasets of 1000 manufactured steady diffusion solutions, generated in the same way as dataset C.
@@
Figure 9: The input structure of an NGO compared to input structure of neural operators from literature like the DeepONet, FNO and VarMiON. Typically, the learnable components of canonical neural operators take as input the input function $\theta$ sampled on a quadrature grid $\mathbf{x}_{q}$, so the size of the input vector is determined by the number of quadrature points $Q$. On the other hand, the NGO system net takes as input inner products of the input function with a basis, which are computed using a quadrature rule with weights $w_{q}$ and points $\mathbf{x}_{q}$. The size of the NGO input is thereby determined by the size of the basis $N$, and independent of the number of used quadrature points $Q$. Therefore, for the same NN input size (in the example illustration, 5 numbers), an NGO can use (arbitrarily) finer quadrature, resulting in higher accuracy on problems that involve fine scale data, where dense quadrature is required.
@@ 4.3.6 NGO-based Preconditioners
Recently, neural operators have gained interest as methods to speed up the numerical solution of PDEs, using discretizations such as finite differences or finite elements. There are multiple ways to achieve this acceleration. One method, used by Zhang et al. [62], Hu et al. [21], and Ackmann et al. [1] among others, constructs a fixed-point iteration scheme for linear systems (derived from PDEs), that alternates between a “standard” smoothing step, such as Jacobi or Gauss-Seidel, and an application of the respective neural operator. Others, including Kopaničáková et al. [kopaničáková2025deeponet] and Xiang et al. [60], use the learned operator as a preconditioner directly, in combination with a linear solver such as Flexible GMRES [52] to allow nonlinear operators as preconditioners for linear systems. However, both approaches require specific choices for the solver or iteration scheme used, which limits their applicability to existing PDE solver software.
An advantage of NGOs is that the fact that they learn an approximation to the Green’s function facilitates the construction of preconditioners for numerical discretizations of the same PDE. Constructing preconditioners from Green’s functions can be done qualitatively, as done by Loghin [36], or numerically using learned Green’s functions as done by Ichimura et al. [25]. In this section, we give an example of using a trained NGO to construct a preconditioner for a finite difference discretization for the diffusion equation. Specifically, a given discretization of a PDE yields a linear system with a matrix $\mathbf{C}[\theta]$, for which we can construct a preconditioner $\mathbf{M}[\theta]$ from the NGO, such that the condition number of $\mathbf{P}=\mathbf{C}\mathbf{M}^{-1}$ is much smaller than that of $\mathbf{C}$.
For this example, we take the steady diffusion problems from the testing data set, formulate a finite difference discretization of the PDE, and use the data NGO trained in Section 4 to construct a preconditioner. For the finite difference discretization, we assume a uniform grid (i.e. $\Delta x=\Delta y\equiv h$) and use a five-point difference scheme to approximate the differential operator:
@@
where $\theta_{i,j}=\theta(x_{i},y_{j})$ and $u_{i,j}$ is the finite-difference approximation to $u(x_{i},x_{j})$. We take $h=\frac{1}{99}$, which produces a discretization with $100$ points in the $y$-direction and $98$ points in the $x$-direction (the Dirichlet boundary conditions imply that no boundary points need to be included). This produces an $9800\times 9800$ linear system $\mathbf{C}\mathbf{u}=\mathbf{b}$.
To precondition this linear system, we use the direct preconditioning (DP) approach described in Kopaničáková et al. [kopaničáková2025deeponet]: the linear system is preconditioned using a multiplicative combination of a smoother and a coarse-level preconditioner, with the latter represented by a neural network. In this work, block-Jacobi is used as a smoother, and the coarse-level preconditioner is given by either a DeepONet or an NGO. To apply a DeepONet or an NGO as a preconditioner for such an $9800\times 9800$ linear system $\mathbf{Cu}=\mathbf{b}$, the model must be able to take as input any vector $\mathbf{v}\in\mathbb{R}^{9800}$ and return an approximation for $\mathbf{C}^{-1}\mathbf{v}\in\mathbb{R}^{9800}$. Since $\mathbf{C}^{-1}\mathbf{v}$ corresponds to the solution of a diffusion equation where the right-hand-side terms correspond to the vector $\mathbf{v}$, this can be done by converting the vector $\mathbf{v}$ to a representation taken by the neural operator, and then applying said operator to obtain an approximation for the solution to the diffusion equation on the desired grid points.
Preconditioning with NGOs
To use an NGO as a coarse-level preconditioner, we note that the action of a Green’s operator can be approximated on the points used by the discretization:
$\hat{u}(\mathbf{x}_{i})$ $\approx\int_{\Omega}\hat{G}[\theta](\mathbf{x}_{i},\mathbf{x}^{\prime})\,f(\mathbf{x}^{\prime})\,d\mathbf{x}^{\prime}\approx\sum_{j=1}^{n}\hat{G}[\theta](\mathbf{x}_{i},\mathbf{x}_{j})\,f(\mathbf{x}_{j})\,w_{j},$ (4.19)
where $\mathbf{x}_{j}$ are the points used by the finite-difference discretization, and $w_{j}$ are associated quadrature weights. NGOs approximate the Green’s function as in (2.8):
$\hat{G}[\theta](\mathbf{x}_{i},\mathbf{x}_{j})$ $\approx\sum_{m,n}\phi_{m}(\mathbf{x}_{i})\mathbf{A}_{mn}[\theta]\phi_{n}(\mathbf{x}_{j}).$ (4.20)
Since the right-hand-side vector $\mathbf{v}$ contains pointwise evaluations of the source term, i.e. $\mathbf{v}_{j}=f(\mathbf{x}_{j})$, we can write the discretized action of the NGO as
$\mathbf{C}^{-1}\mathbf{v}\approx\mathbf{P}_{\text{NGO}}(\mathbf{v}):=\mathbf{PA[\theta]R}\mathbf{v},\text{ where }\mathbf{P}_{ij}=\phi_{j}(\mathbf{x}_{i}),~\mathbf{R}_{ij}=\phi_{i}(\mathbf{x}_{j})w_{j}.$ (4.21)
Therefore, the action of the NGO as a preconditioner as represented by the sequence of operations on $\mathbf{v}$ in (4.21) may be conceived of as a coarse-level preconditioner in the following manner:
@@
    Computing the inner product of $\mathbf{v}$ with the NGO basis functions, denoted by the operator $\mathbf{R}_{ij}$, represents a restriction operator.
@@
    The matrix $\mathbf{A}[\theta]$ computed by the NGO represents a coarse-scale solution operator.